When I get to the last portion of the problem I get stuck on this.
5/9 * 9/5 < 5/9 (1542.4)
The answer comes out to
C < 856.9 but how did they get it?

Found 2 solutions by ilana, solver91311:
Answer by ilana(307)

You left out a piece of the inequality- the variable C.
It looks like the inequality was originally 9/5 * C < 154.2.
Then, to isolate the variable C, you multiply each side of the inequality by 5/9. Since this is a positive number, the inequality symbol stays the same.
Now you have (almost) what you wrote.
5/9 * 9/5 * C < 5/9 * 1542.4
Then multiply on the left side, 5/9 * 9/5 = 1, so 5/9 * 9/5 * C = C.
So the inequality is C < 5/9 * 1542.4
Now multiply on the left side, 5/9 * 1542.4 = (approximately) 856.9.
So the inequality is C < 856.9

Answer by solver91311(24713)





which, rounded to the nearest tenth is 856.9.

However, since there is no variable "C" in your original inequality, I don't know how to get to



from



The best I can do is the triviality:



I suspect you have made an error in an earlier step. Share the original problem please. Share it on the board so everyone can see it.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism